Second LBC Tech Night

Data on:/Repository/20091014_Tech
Observers: Testa, V, Pedichini F. for LBC team, Pedani, M. for LBTO

Cirrus clouds T_air = 5.0 RH~65% Wind up to 15 m/s

LBC Team

Change Blue shutter ok

noise on red

01:30 - SkyFlatTest_UY

01:35 - SkyFlat_UY 5s many stars, end flats and do some IA

01:38 - dofpia B: 014050 - -14278 0 0 0 0 0 0 2.83 5.63 R: 014051 - 5789 -810 62 -2136 -896 0 0 1.57 2.52

Blue is wrong!

B: 014304 - error, undetected coma R: 014300 - -1235 -157 -432 -27 73 0 0 1.76 0.51

Blue: Z4=8000 Z7=-2000

01:46 - dofpia

B: 014729 - -13000 white spot in the center of the pupil.... R: 014725 - 622 616 -195 -52 104 0 0 1.63 0.32

B: error R: 93 -85 181 98 -43 0 0 1.55 0.12

clear active optics in blue Z7 = -2500 Z8 = +1000

01:54 - dofpia

B: 960 565 -613 -1632 -1082 -196 100 0.00 1.00 R: 708 32 149 -36 87 -253 36 1.35 0.34

B: -1393 -891 86 -100 -244 0 0 1.78 0.63 R: -285 252 -196 222 -10 -22 41 1.29 0.19

B: -301 -105 -338 47 35 0 0 1.89 0.18 R: -41 -144 52 -81 20 76 27 1.25 0.11

02:07 Slew to ACT0304 to adjust Pointing

Old CA= -28.95 Old IE= -23.17

2009/10/14 02:11:46.965838 CAMERA SHUTTER timeout error on communication >XPS :?< [src/camera/camera.c:1458] 2009/10/14 02:11:46.965930 CAMERA SHUTTER shutter communication problem: telemetry failure correction set to average value

No Red image. Shutter failed.

lbcr.20091014.021322.fits

New CA=-57.2 New IE=-40.5

lbcr.20091014.022035.fits

New CA=-51.6 New IA=-38.9

lbcr.20091014.022354.fits AZ=176 now

New CA=-47.3 New IE=-39.1

Note the drift of CA with Azimuth Ok we choose these values for CA,IE

02:30 DOFPIA again

R: -1061 372 -133 342 -148 153 -40 0.78 0.48 B: 308 423 -102 23 -22 185 98 1.32 0.23

R: 267 2 248 -120 -110 -48 -27 0.87 0.17 B: -325 -264 -35 134 71 105 7 1.29 0.19

02:34 Test Rotator Trajectory

lbcb.20091014.023522.fits 180s V-band EL=+85 fwhm=0.9"
lbcr.20091014.023516.fits 180s V-band EL=+85 fwhm=0.9"

Stars pretty well rounded in 180s exposures. Rotator trajectory Calculation looks OK.

IQ Script Results:

### Seeing data for LBC, UT=20091014
### Run time: Tue Oct 13 22:58:25 MST 2009
### UT(hr) FWswl IMAGE CAM FILT Texp AIRM FWHM Ellip Np
2.589 1.01 023522 B Vb 180.24 1.004 4.48 0.082 6195
2.588 1.01 023516 R Vb 180.23 1.004 4.48 0.052 4403

Blue is a bit more elongated but it would pass the IQ threshold of 0.1

02:45 Test Pointing hysteresis

EL=+84

lbcb.200910014.024703.fits (x=1223.61 y=2232.68)
lbcr.200910014.024658.fits (x=1137.79 y=2572.64)

Then Slewing manually at EL=+20

Now again at EL=+84

lbcb.200910014.025038.fits (x=1248.24 y=2254.77)
lbcr.200910014.025033.fits (x=1140.0 y=2576.6)

NO POINTING HYSTERESIS FOUND with LBCR (~1" movement, compatible with no guiding for a while) For LBCB we FIND: 5.5" at X and 4.9" at Y movement which translates into: ~0.5" in AZ and ~+7.0" in EL.

03:10 Tests Aberrations for Andrew (We introduce aberrations and we take intra+ext

03:11 Collimate EL=+78

We put Z4=-2000nm

RB_rVfastextra B: 031708 R: 031705

Collimate RB_rVfastintra B: 032519 R: 032519

We put Z4=+2000nm

Recollimate

We put Z4=-1500nm

RB_rVfastextra B 33023 R 33019

RB_rVfastintra B 33150 R 33144

We put Z4=+1500nm Recollimate

We put Z4=-1000nm

RB_rVfastextra B 033834 R 033830

RB_rVfastintra B 034001 R 033955

We put Z4=+1000nm Recollimate

We put Z4=-500nm

RB_rVfastextra B 034514 R 034510

RB_rVfastintra B 034639 R 034631

We put Z4=+500nm Recollimate

We leave now Z4=0.0

RB_rVfastextra B 035037 R 035033

RB_rVfastintra B 035202 R 035153

We put Z4=+500nm

RB_rVfastextra B 035502 R 035455

RB_rVfastintra B 035628 R 035626

We put Z4=-500nm Recollimate

We put Z4=+1000nm

RB_rVfastextra B 040113 R 040108

RB_rVfastintra B 040253 R 040240

We put Z4=-1000nm Recollimating

We put Z4=+1500nm

RB_rVfastextra B 040824 R 040820

RB_rVfastintra B 040950 R 040941

We put Z4=-1500nm Recollimate

We put Z4=+2000nm

RB_rVfastextra B 041514 R 041511

RB_rVfastintra B 041653 R 041639

End tests; Z11 and Z22 sequences TBD yet.

04:18 Leave telescope to LBC team for IQ analysis work.

-- MarcoPedani - 14 Oct 2009
I Attachment Action Size Date Who Comment
daq.20091014.gifgif daq.20091014.gif manage 45 K 14 Oct 2009 - 23:38 JohnHill Plot of Telescope and Mirror Temperatures
Topic revision: r5 - 15 Oct 2009, MarcoPedani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback